Shift report for DAY 17/8/99

Elliott Cheu (elliott@physics.arizona.edu)
Tue, 17 Aug 1999 13:45:01 -0700

Shift Summary for DAY shift, 17/08/99

On Shift: ECC, BW

Hours of beam available at all intensities: 1.5
Hours of beam available at nominal intensity: 1.5
Hours of E832 data (special runs not included): 1.5
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Machine studies.

========================================================================
Runs ended this shift: 14242
Run 14242 - 12.1M events - BEAM_832

Tapes brought to FCC this shift: None

========================================================================
Accesses during shift:
- Get DC samples
- Replace CsI HV cable (1137)
- Move DLTs from loading dock
- Measure DC Z positions and improve pump oil readability
- Fix various CsI channels (226 216)

========================================================================
Overview of shift:
E832 data taking and experiment improvements while no beam.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14242 180 7:30 SEM > 7.5e12.
8:25 Beam goes away.
8:45 New DOT installed to improve tape "appending".
8:50 Interlocks dropped.
9:10 Tapes moved into KTeV Hall.
10:40 HCC Dynode (1799) examined. May have been
loose at rib. Looks all right now.
10:44 Hot wire in 1Y-4. Fixed after reseating cables.
11:30 Laser light boosted by 8-9K pin diode counts.
12:00 Spare HV cable installed for CsI 1137.
Channels 226, 216 now seem to be flaky.
13:30 Fixed trigger bit 16.
BW labelled tapes.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Fix CsI channels.